# StarEmbed: Benchmarking Time Series Foundation Models on Astronomical Observations of Variable Stars

### v1

@@ 7 Acknowledgement
The first three authors, Weijian Li, Hong-Yu Chen, and Qinjie Lin, contributed equally to this work. The remaining authors provided additional support and contributions.
@@ Appendix A Full List of Hand-crafted Features
We select hand-crafted features from the libraries of established software packages: FATS (I. Nun et al., 2015) and light_curve (K. L. Malanchev et al., 2021a). Each feature described here is computed for each passband individually and the embeddings are formed by concatenating the feature lists of the $g$ and $r$ embeddings. Tables 7 and 8 show the full list of features and descriptions from FATS and light_curve, respectively.
@@ B.1 Eclipsing Binaries (EW, EA)
EW-type eclipsing binaries (also called W Ursae Majoris-type after the original EW system) are contact binaries. In this case, the two stars, typically dwarf stars, share a common outer envelope which entirely encapsulates them. This common envelope allows for the exchange of mass and energy between the pair, equilibrating their temperature. Their light curves exhibit constant and smooth variations where the dips from either star being eclipsed are of similar or identical depth. EW-type variables typically have short periods ($0.2\lesssim P\textrm{ [d]}\lesssim 0.5$), with a notably unsolved period cut-off at $\sim 0.2$ days (S. M. Rucinski, 2007; A. J. Drake et al., 2014a). These systems are astrophysically valuable, in part, because they are expected to emit gravitational waves due to their tight orbits, and they also have the possibility of merging and triggering transient events (R. Tylenda et al., 2011).
@@ B.4 Long Period Variables (LPVs)
LPVs are giant stars exhibiting pulsations with periods of $3\lesssim P\textrm{ [d]}\lesssim 1000$. They include multiple subtypes each with different period-luminosity relations but we, as in A. J. Drake et al. (2014a), consider these a single class. They pulsate with a similar mechanism to the RR Lyrae but have dramatically larger radii. Their outer layers are not very tightly bound to the rest of the star, which can lead to the star expunging mass and polluting the surrounding environment with gas. Thus, studying LPVs provides insights into the cycles of gas into and out of stars. Their period-luminosity relations also allow LPVs to act as distance measures.
@@ B.10 Hump variables
The Hump class is used as a catch-all for the small amount of periodic variables which A. J. Drake et al. (2014a) were unable to classify into any other known classes but do show clear periodic variability. Some of these objects exhibit vaguely sawtooth-like variability, like what is seen in RRab stars, but others have smoother variability.
@@ B.12 EA with unknown period (EAup)
EAup are EA-type stars where A. J. Drake et al. (2014a) were unable to determine their periods for any reason.
@@ Appendix C Astromer Performance and Astromer-1 Embedding Quality
We provide detailed analysis and provide experiments on the issue of the poor performance of Astromer-1. First, Astromer-1 needs further finetuning on the dataset of the downstream task to achieve good performance on the variable star classification task, according to C. Donoso-Oliva et al. (2023). This is evidenced in their Fig. 11 (a), which shows a clear increase of F1 score from  0.25 to  0.6 when finetuned on 20 to 500 variable stars per class of the MACHO dataset. A similar trend is observed for other datasets, including OGLE-III and ATLAS. Astromer-2 (C. Donoso-Oliva et al., 2025) also shows improvement with finetuning, though its performance starts higher (around 0.65) even with just 20 samples per class. Please refer to (C. Donoso-Oliva et al., 2025, Figs. 11 and 12) for the details. These results indicate that Astromer’s low zero-shot performance in our benchmark is expected, as we intentionally evaluate the pretrained checkpoints without any task-specific tuning.
@@
Appendix D Additional Experiments Results on Confusion Matrix
@@
To provide detailed information about classification performance across different variable star classes, we report the confusion matrices for all embeddings and the hand-crafted features in this section. For the random forest and MLP classifiers, the per-class accuracies are obtained by averaging over the results of three runs.

### v2

@@ VII Acknowledgement
The first three authors, Weijian Li, Hong-Yu Chen, and Nabeel Rehemtulla, contributed equally to this work. The remaining authors provided additional and substantial contributions.
@@ Appendix A Full List of Hand-crafted Features
We select hand-crafted features from the libraries of established software packages: FATS (Nun et al., 2015) and light_curve (Malanchev et al., 2021a). Each feature described here is computed for each passband individually and the embeddings are formed by concatenating the feature lists of the $g$ and $r$ embeddings. Tables 7 and 8 show the full list of features and descriptions from FATS and light_curve, respectively.
@@ B.1 Eclipsing Binaries (EW, EA)
EW-type eclipsing binaries (also called W Ursae Majoris-type after the original EW system) are contact binaries. In this case, the two stars, typically dwarf stars, share a common outer envelope which entirely encapsulates them. This common envelope allows for the exchange of mass and energy between the pair, equilibrating their temperature. Their light curves exhibit constant and smooth variations where the dips from either star being eclipsed are of similar or identical depth. EW-type variables typically have short periods ($0.2\lesssim P\textrm{ [d]}\lesssim 0.5$), with a notably unsolved period cut-off at $\sim 0.2$ days (Rucinski, 2007; Drake et al., 2014a). These systems are astrophysically valuable, in part, because they are expected to emit gravitational waves due to their tight orbits, and they also have the possibility of merging and triggering transient events (Tylenda et al., 2011).
@@ B.4 Long Period Variables (LPVs)
LPVs are giant stars exhibiting pulsations with periods of $3\lesssim P\textrm{ [d]}\lesssim 1000$. They include multiple subtypes each with different period-luminosity relations but we, as in Drake et al. (2014a), consider these a single class. They pulsate with a similar mechanism to the RR Lyrae but have dramatically larger radii. Their outer layers are not very tightly bound to the rest of the star, which can lead to the star expunging mass and polluting the surrounding environment with gas. Thus, studying LPVs provides insights into the cycles of gas into and out of stars. Their period-luminosity relations also allow LPVs to act as distance measures.
@@ B.10 Hump variables
The Hump class is used as a catch-all for the small amount of periodic variables which Drake et al. (2014a) were unable to classify into any other known classes but do show clear periodic variability. Some of these objects exhibit vaguely sawtooth-like variability, like what is seen in RRab stars, but others have smoother variability.
@@ B.12 EA with unknown period (EAup)
EAup are EA-type stars where Drake et al. (2014a) were unable to determine their periods for any reason.
@@ Appendix C Astromer Performance and Astromer-1 Embedding Quality
We provide detailed analysis and provide experiments on the issue of the poor performance of Astromer-1. First, Astromer-1 needs further finetuning on the dataset of the downstream task to achieve good performance on the variable star classification task, according to Donoso-Oliva et al. (2023). This is evidenced in their Fig. 11 (a), which shows a clear increase of F1 score from  0.25 to  0.6 when finetuned on 20 to 500 variable stars per class of the MACHO dataset. A similar trend is observed for other datasets, including OGLE-III and ATLAS. Astromer-2 (Donoso-Oliva et al., 2025) also shows improvement with finetuning, though its performance starts higher (around 0.65) even with just 20 samples per class. Please refer to (Donoso-Oliva et al., 2025, Figs. 11 and 12) for the details. These results indicate that Astromer’s low zero-shot performance in our benchmark is expected, as we intentionally evaluate the pretrained checkpoints without any task-specific tuning.
@@
Appendix D Feature Importance Analysis for Hand-crafted Features
In this appendix we investigate the feature importance of the individual hand-crafted features to investigate the source of their excellent performance. We perform this analysis on the random forest trained for supervised classification with the hand-crafted features because it offers built-in functionality for quantitatively measuring the importance of each feature. The feature importance is computed using the mean decrease in impurity. In this framework, each time a feature is used to split a node, the associated reduction in the Gini impurity metric is recorded and weighted by the number of samples reaching that node. These weighted impurity decreases are then summed over all nodes and all trees in the ensemble, and finally normalized to yield a global, unit-scaled measure of how strongly each feature contributes to improving class separation within the forest.
Figure 7: Top six most important hand-crafted features for random forest classification. The period of variability computed for either passband rank the highest. Other important metrics also include the skewness of the distribution of magnitudes, the median absolute deviation, and the magnitude percentage ratio for the 40th and 5th percentiles.
Figure 7 shows the results of our feature importance analysis. We find that the period inferred from the $g$ and $r$ bands of the time series are the most important features driving the performance of the hand-crafted features. This aligns with expectations as numerous other studies performing variable star classification with tree-based classifiers come to a similar conclusion Richards et al. (2011); Dubath et al. (2011); Kim and Bailer-Jones (2016). We also find that the skewness, median absolute deviation, and the magnitude percentage ratio also rank very highly. These features encode information about the spread of the magnitude measurements, indirectly representing the morphology of the periodicity of the star. All features are described in Appendix A.
Appendix E Preliminary Results of Fine-Tuning TSFMs on Astrophysics Data
Our goal in this work is to answer a very specific first-order question: how useful are existing, large TSFMs without any model adaptation for real, irregular astrophysical light curves? This is fundamentally a question of generalization capabilities, motivated by strong claims made by the latest-generation TSFMs (Chronos, Chronos-bolt, Moirai). To this end, we explicitly treat these TSFMs as frozen embedding models and do not fine-tune them on our data. This “zero-shot” setting on new, out-of-domain data is precisely where generalization abilities can be tested.
Methodologically, we do go beyond pure zero-shot prediction by training lightweight supervised heads (k-NN, linear probe, RF, MLP) on top of frozen embeddings (see Sec. IV.2). This isolates the question we care about: do the representations produced by TSFMs, without architecture- or domain-specific fine-tuning, perform competitively with conceptually very simple baselines? Our results show that they do: TSFMs approach or surpass these baselines for clustering and OOD detection and come close for supervised classification, despite never being trained on astronomical data. This alone is a strong and surprising conclusion.
To better understand the potential of adaptation, we also conduct a preliminary, and deliberately simple, fine-tuning experiment on Moirai-small. We fine-tune the model on about 1M ZTF light curves for about 24 hours using the same masked token prediction task as in the pre-training. As shown by Figure˜8, the training loss decrease from 1.25 to 1.15 within the first 20k steps but then enter a regime of strong fluctuations around a saturated plateau over the next 180k steps, without clear signs of further improvement. This is the evidence that straightforward, off-the-shelf fine-tuning is not trivial and does not immediately yield clear gains over the zero-shot setting. This result in aligned with our interpretations on the shortcomings of the TSFMs: there exists a significant gap between the irregular, sparse astrophysical light curves and the regular time series on which these TSFMs were originally pre-trained. Not treating the irregular sampling of the data destroys much of the important information it holds, see Fig. 5. Fine-tuning these TSFMs on astronomical time series is a non-trivial, model- and domain-design problem in its own right.
Exploring one/few-shot model adaptation–e.g., careful parameter-efficient fine-tuning (PEFT) of selected layers under small labeled budgets, with appropriate regularization and class balancing–would therefore require substantial additional design choices: which layers to adapt, which PEFT strategy, what label budgets, how to avoid overfitting on common classes, etc. This constitutes a large and interesting future research track for follow-up work of our benchmark.
Figure 8: Training loss of fine-tuning Moirai-small on 1 million of ZTF variable star light curves. The train loss saturates quickly after a small improvement and fluctuates around a plateau. Straightforward, off-the-shelf fine-tuning of TSFMs to light curve data requires non-trivial modifications to the model architecture and/or the standard fine-tuning practices and does not immediately yield clear gains over the zero-shot setting.
Appendix F Period Regression Results
As we discuss in Section VI and Appendix D, our hypothesis on the fact that TSFMs underperform hand-crafted feature is due to the failure of capturing the true period of the light curve. Thus, we have added period regression as an additional task in our benchmark.The result is in Table 9.
Model RMSE(d) R2
MOIRAI $26.355\pm 0.022$ $0.295\pm 0.001$
Chronos-bolt $23.414\pm 0.062$ $0.443\pm 0.002$
Chronos $23.178\pm 0.076$ $0.455\pm 0.003$
Astromer-1 $33.272\pm 0.015$ $0.005\pm 0.0009$
Random-emb $33.452\pm 0.016$ $0.005\pm 0.0009$
Table 9: Period regression on the model embedding. While differences do exist in the regression performance of these embeddings, it is important to note that they all perform signficantly worse than Lomb Scargle periodogram which is a reliable, albeit computational expensive, method used for period finding in Astronomy. Note that we exclude hand-crafted features since they already contain the period.
Appendix G Additional Experiments Results on Confusion Matrix
@@
To provide detailed information about classification performance across different variable star classes, we report the confusion matrices for all embeddings and the hand-crafted features in this section. For the random forest and MLP classifiers, the per-class accuracies are obtained by averaging over the results of $10$ runs.